Ammonia is an important compound and useful to industries such as agricultural, pharmaceuticals, textile, plastics, petroleum, mining, and commercial. In the ammonia production, Haber process is utilized and it is shown as, N2(g) + 3H2(g) ➡️ 2NH3(g).

How many grams of ammonia will be produced if there are 500.0 moles of nitrogen gas?
How many grams of hydrogen gas are needed to produce 10,000.0 g of ammonia?
1
Expert's answer
2020-11-23T06:56:31-0500
  1. m(NH3)=n(N2)*2Mr(NH3) = 500*2*17= 17000 g
  2. m(H2)=m(NH3)*3Mr(H2)/2Mr(NH3)= 10000*3*6/(2*17)=5294.12 g